Fermi National Laboratory
 

Accelerator Update for May 26, 2006 - May 30, 2006

 

Friday May 26

The day shift began with beam to the Booster, Main Injector (MI), and Recycler, and with experts accessing the Tevatron (TeV) to investigate a ground fault.

Operations reported at 9:11 AM that vacuum technicians were checking the vacuum at TeV sector B1, and that experts were in the Transfer hall to work on the Tevatron Electron Lens (TEL).

Antiproton Source (Pbar) experts began trying to establish 8 GeV reverse protons at 10:23 AM.

A MI expert began a slip stacking study at 10:24 AM.

Everyone had come out of TeV access by 10:57 AM.

Operations reported the TeV was given permission to power the bus; at 3:40 PM there was current on the TeV bus.

Booster studies ended at 5:21 PM.

Operations established a 980 GeV dry store (no beam) at 10:12 PM. The ramp dumped about three minutes later due to a transductor failure.

Operations reestablished a 980 GeV dry store at 11:36 PM.

Saturday May 27

The midnight shift began with Booster, MI, Pbar, and Recycler studies, and with TeV power supply testing.

At 2:26 AM, the TeV suffered a quench at sector E1. Cryo system experts investigated and reported at 5:30 AM that the E1 wet engine gas bottle was found valved out.

Operations reported at 6:06 AM that the TeV sector E1 cryo system had recovered.

Operations reestablished a 980 GeV dry store at 6:47 AM.

A Pbar expert reported at 8:51 AM that Pbar had established a circulating in the Accumulator.

Recycler experts began making emittance measurements at 4:11 PM.

Pbar experts established beam to the Debuncher at 4:16 PM, but at a very poor efficiency.

Operations established a 980 GeV dry store at 6:05 PM.

Sunday May 28

The midnight shift began with Booster and Pbar conducting studies, with the TeV at 980 GeV, and with the TeV B11 separator baking.

At 8 AM, Booster experts began a four-hour 400 MeV destructive beam study.

At 2:18 PM, a MI enclosure door at MI-20 that hadn’t latched properly blew open due to the winds, dropping the MI interlocks, and causing Recycler to abort its protons. Everything was back on line by 3:41 PM.

Monday May 29

The midnight shift began with Pbar and Recycler conducting studies, and with the TeV ramping.

Pbar experts accessed their rings at 9:37 AM. They came out of access about forty-five minutes later.

Operations reported at 2:30 PM that the TeV bakeout was finished and that experts were going to access the TeV.

At 4:22 PM, Operations reported that lightning had caused eight Booster RF stations, the MI PSLOOP, Booster vacuum valves and chopper, and some 4000 MeV quadrupole power supplies to trip off.

Operations established a dry squeeze at 10:43 PM.

Tuesday May 30

The midnight shift began with Recycler and Pbar conducting studies, and with the TeV in a dry squeeze.

At 1:24 AM, Operations reported a raccoon attack on the Linac gallery. It seemed to be a coordinated effort. Fortunately, by 1:53 AM, a joint force of operators and Pbar experts managed to drive the raccoons out of their hastily made fortifications. Then at 4:18 AM, the raccoons made what some thought to be a counter attack on the Division Headquarters, but others believed it to be only a simple reconnaissance incursion. No raccoons were either injured or captured during these encounters. Operator losses were low.

Operations prepared Pbar for an access at 5:44 AM.

Booster experts began a study period at 7:54 AM.

The Plans for Today

The plans are to continue with startup studies. There will be TeV safety system tests tonight. The Collider Detector and Fermilab and D-Zero will be interlocked for these tests. Experts may access the TeV tomorrow. The Run Coordinator hopes to have beam in the TeV by Thursday.

Machine Reports

Linac

LRF1 and LRF5 have problems that are being investigated.

Booster

Experts tuned up their machine. They have a problem with the vertical IPM that needs further investigations.

MI

The MI-8 line is okay. Experts need to set the 120 GeV tune for NuMI. Experts also want to conduct an aperture scan.

Pbar

Experts discovered that they have a broken quadrupole shunt. They believe they are a couple days away from making antiprotons.

TeV

Experts believed that their ground fault was due to water on the spool piece at sector A48. The B11 separator bake was successful. They continue to condition separators.

Recycler

Experts held a long store of protons. The Electron Cooling SF6 cooling skid still isn’t working. Experts have discovered that they can run the Pelletron for about eight hours with only exterior cooling, and then they need to let it cool for sixteen hours.

Switchyard 120

Waiting for a B2 dipole magnet to be found to replace the bad one in the P3 line.

MiniBooNE

Experts are ready for beam. Unfortunately, the Booster isn’t ready to deliver the intensity MiniBooNE needs.

NuMI

The water mitigation work is finishing up. The chillers were run all weekend. Experts want a search and secure today and power supply testing tonight. They would like some beam tomorrow.

Collider Detector at Fermilab

COT techs found and fixed a bad ground that was causing their noise problems. They have closed up their plug, but not bolted it in yet. They will be ready for an access after 4 PM today.

D-Zero

Experts continue to check out their detector. They expect the survey crew to be finished this afternoon. They will be ready for an access after 4 PM today.


Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now.

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.

last modified 05/30/2006    email Fermilab